OpenStudy (anonymous):

g(x)=x^2-10x-3 g(x+5) when i get this far on x^2+10x+25-10x-50-3 next step is to order terms before any math is completed?

OpenStudy (anonymous):

IF I am not mistaken, combine like-terms is next.....

OpenStudy (zzr0ck3r):

why did you repost this?

OpenStudy (anonymous):

cus u left befor clarity

OpenStudy (zzr0ck3r):

did you ask a question?

OpenStudy (anonymous):

cus u left before clarity

OpenStudy (zzr0ck3r):

I did not leave I looking at others, I can multi task

OpenStudy (zzr0ck3r):

again, did you ask a question?

OpenStudy (anonymous):

yes, the one that i reposted

OpenStudy (zzr0ck3r):

lol, ok next time... if I answer and then leave, ask a question...

OpenStudy (zzr0ck3r):

don't repost

OpenStudy (zzr0ck3r):

there is no point...

OpenStudy (anonymous):

i see now, but i just seen everyone abandoned my question rather or it was intended to be temporary, i didnt know. i guess a little impatient, ive been online and hadnt got that far!

OpenStudy (anonymous):

i dont come for yall to do my ork. i want to learn, not get just get good grade on homework

OpenStudy (anonymous):

work*

OpenStudy (zzr0ck3r):

x^2+10x+25-10x-50-3 we regroup so the like terms are next to each other x^2+10x-10x+25-50-3 +10x-10x = 0 so x^2+25-50-3 25-50-3 = -28 so we have x^2-28

OpenStudy (zzr0ck3r):

I understand

OpenStudy (anonymous):

ok go multi task while i look at what u have before me and ill message u if i need you, deal?

OpenStudy (zzr0ck3r):

Im just saying, just because we leave the question does not mean we are not coming back. And if you want me back then do this @zzr0ck3r

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

don't even msg me just ask the question, its posts and tells me when there are more comments on any question I have worked on, so I always try and look to see what they are:)

OpenStudy (anonymous):

but go head and help someone and imma look at what u put up so i can try to understand

OpenStudy (zzr0ck3r):

let me know when you want to talk about g(-x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

bye for now

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

yo

OpenStudy (zzr0ck3r):

I have to go pick my brother up at the airport, so make this quick or wait about 30 mins...

OpenStudy (anonymous):

g(-x)=x^2=10x-3?

OpenStudy (anonymous):

remember -x

OpenStudy (zzr0ck3r):

should that = be a +?

OpenStudy (zzr0ck3r):

if so then yes:)

OpenStudy (anonymous):

where a +

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

g(-x)=x^2+10x-3

OpenStudy (zzr0ck3r):

x^2+10x-3?

OpenStudy (zzr0ck3r):

yeah

OpenStudy (anonymous):

have a safe trip

OpenStudy (zzr0ck3r):

nice work:) most people would have said -x^2

OpenStudy (anonymous):

thank you